### v2

# Unleashing HyDRa: Hybrid Fusion, Depth Consistency and Radar for Unified 3D Perception

Abstract
Low-cost, vision-centric 3D perception systems for autonomous driving have made significant progress in recent years, narrowing the gap to expensive LiDAR-based methods. The primary challenge in becoming a fully reliable alternative lies in robust depth prediction capabilities, as camera-based systems struggle with long detection ranges and adverse lighting and weather conditions. In this work, we introduce HyDRa , a novel camera-radar fusion architecture for diverse 3D perception tasks. Building upon the principles of dense BEV (Bird’s Eye View)-based architectures, HyDRa  introduces a hybrid fusion approach to combine the strengths of complementary camera and radar features in two distinct representation spaces. Our Height Association Transformer module leverages radar features already in the perspective view to produce more robust and accurate depth predictions. In the BEV, we refine the initial sparse representation by a Radar-weighted Depth Consistency. HyDRa  achieves a new state-of-the-art for camera-radar fusion of 64.2 NDS ($+1.8$) and 58.4 AMOTA (+1.5) on the public nuScenes dataset. Moreover, our new semantically rich and spatially accurate BEV features can be directly converted into a powerful occupancy representation, beating all previous camera-based methods on the Occ3D benchmark by an impressive 3.7 mIoU. Code and models are available at https://github.com/phi-wol/hydra.
1 Introduction
Reliable and affordable 3D perception is an important cornerstone for safe and efficient operation of autonomous vehicles, enabling agents to navigate dynamic environments and complex scenarios. In this rapidly evolving field, deep learning-based Camera-LiDAR fusion has emerged as the de-facto standard for 3D reconstruction from multiple sensors, with a proven track record in 3D object detection [77, 45, 37], semantic segmentation [45, 12, 49], and object tracking [8, 37].
While delivering highly accurate geometry sensing, LiDAR sensors still pose a high-cost barrier for large-scale deployment. In order to democratize access to autonomous driving systems, LiDAR-free, and vision-centric systems have gained significant traction in recent years and months. Especially with the advent of Bird’s Eye View (BEV) based architectures [56], camera-only methods have achieved remarkable progress [33, 67, 62]. Cameras provide dense and high-resolution semantic information, theoretically enabling the detection of objects at far distances. However, the main practical challenge lies in preserving the spatial information after 2D projection, leading to unstable depth estimation and localization errors [33, 55]. Another difficulty of passive sensors is posed by dynamic objects that are occluded, yet remain relevant for safe decision making.
Still underrepresented in this research domain is the incorporation of radar sensors. Especially camera-radar fusion has been neglected by the academic community [12]. As low-frequency active sensors based on time-of-flight principles, they exhibit a remarkable resilience to adverse weather and lighting conditions, offering metric measurements for perception ranges up to 300 meters [78]. Furthermore, by leveraging the Doppler effect, radar reflections can yield valuable velocity information about moving and occluded objects, underscoring the potential benefits as a camera-complementary sensor. Combining the two sensor types is key to unlocking the full potential of accessible 3D perception systems [60].
Nevertheless, the task of integrating both sensor types into a coherent and robust perception system is a challenging, non-trivial task. Effectively associating radar detections with camera features is complicated by the unique characteristics of the radar data. The absence of elevation information, the sparse nature of the radar returns, the presence of noise from multi-path reflections, the occurrence of false positives, and the overall high measurement uncertainty make radar data difficult to process and associate with camera features [23, 60]. Consequently, the direct application of state-of-the-art LiDAR-like architectures on inherently sparse radar data has proven to be ineffective [51]. Despite being in point-cloud form, Yan et al. highlight that CMT [74], among the current leading methods on the nuScenes 3D object detection task, degrades in performance when incorporating the radar data into their pipeline. For this reason, we see the need for a significant shift in fusion paradigms when adopting radar. Traditionally, fusion approaches have been heavily biased towards LiDAR, with the bulk of the network’s capacity devoted to processing LiDAR point clouds, and camera data merely serving as a supplementary source of semantic information [65, 31, 19]. Leaving the LiDAR sensor out of the equation, new approaches significantly prioritize extracting dense semantic information from camera feeds and its projection into the BEV grid. The methodology relies heavily on advanced 2D feature extraction techniques [26, 43, 44] along with efficient view transformation and lifting modules [17, 34, 36]. In this new configuration, radar data assumes a role akin to the previous camera-based methods, enhancing the primary modality with sparse metric information and becoming a guidance module [12, 23].
Modern practices have focused on adaptive fusion in the BEV, reducing spatial misalignment during feature aggregation [45, 23]. Despite these improvements, they still struggle with robust depth estimation and rely partly on monocular depth cues for feature transformation [33]. Existing techniques, while innovative, do not fully capitalize on radar’s potential for unified robust depth sensing, often resulting in a dual-stage projection process where each modality stage relies on its own, potentially inconsistent, depth estimation. We argue that to fully unlock the potential, we must move the fusion stage even earlier. In this work, we propose HyDRa, a state-of-the-art camera-radar fusion architecture well equipped to tackle a variety of 3D perception tasks. Our contributions and important stepstones to enable LiDAR-free autonomous vehicles and advanced driver assistance systems are the following:
Refer to captionRefer to caption
Figure 1: Overall Architecture of HyDRa: Modality-specific backbones extract multi-view image and radar pillar features. The features are fused in two representation spaces: Perspective View and BEV-Space. 1. The radar features are associated with the image features by the Height Association Transformer. With the resulting radar-informed dense depth distribution, the forward projection module generates a sparse BEV representation. 2: The splatted semantic BEV features and radar-BEV features are concatenated and fused by a SE block. 3: A depth-aware backward projection refines this representation, guided by radar attention weights before being distributed to task-specific heads.
  1. 1.
    We introduce our Height Association Transformer (HAT) module and address the limitations of previous BEV-generating depth networks. By creating unified geometry-aware features in the perspective view, we significantly reduce the translation error of existing methods.
  2. 2.
    We upgrade the concept backprojection-as-refinement and establish a Radar-weighted Depth Consistency (RDC) for enhancing sparse fusion features in the BEV, tackling the issue of misaligned or unassociated features and occluded objects.
  3. 3.
    We extensively ablate our design choices and show that each of our contributions significantly improves the performance of our model on the nuScenes detection and tracking challenges [2], but the synergy of all components is crucial for achieving best-in-class results.
  4. 4.
    HyDRa  establishes a pioneering model for camera-radar based 3D semantic occupancy prediction on the Occ3D benchmark [61], pushing the upper bound of state-of-the-art low-cost vision-centric methods.
2 Related Work
2.1 Camera-based Architectures
Dense BEV-based Architectures. Shifting from perspective-view 3D object detection [50, 86, 59, 68, 69] to Birds-Eye-View (BEV) based perception [56, 57], the BEV-based architectures have become the new standard for 3D object detection and tracking [28]. With the introduction of differentiable lifting methods, the image-based output of a standard 2D convolutional neural networks [13, 44] or a transformer-based feature extractor [9, 43] is transformed into the more downstream task-friendly BEV ground plane representation [86, 15]. The BEVDet series [18, 16, 17] and following work [45, 33] have raised the bar for efficient and powerful forward projection pipeline, predicting an explicit, dense, and pixel-wise depth distribution and pooling the camera-frustum features in the BEV space. BEVDepth [33], SoloFusion [55], and BEVStereo [32] focused on localization by leveraging LiDAR supervision during training, investigating long-term 4D fusion and temporal stereo matching for better depth estimation. BEVFormer [34, 75] has shown that inverse modeling of the view transformation[58] is a valid alternative to the forward projection pipeline; by back-projecting a set of pre-defined BEV grid queries. Spatial deformable cross attention [64, 88] implicitly learns to model the 3D relations by interacting with regions of interest across multiple image feature maps.
FB-BEV [36] combines both approaches into the forward-backward projection paradigm, modeling a more precise projection relationship. It addresses the sparsity and inaccuracy inherent in both projection types by ensuring a Depth Consistency between the two modules. We will build upon this promising concept and extend it to the multi-modal domain by introducing a novel Radar-weighted Depth Consistency and incorporating radar in complementary representation spaces to enforce a more accurate depth estimation.
Sparse BEV-based Architectures. Instead of modeling the complete surrounding, a concurrent line of research has established the modeling of the scene as a sparse set of object queries [4, 27, 82], decoding 3D proposals via implicit attention [41, 70, 42, 66]. In addition to long-term temporal fusion [55], StreamPETR introduces an object-level temporal mechanism and query denoising [27, 82] to establish an efficient and powerful detection and tracking paradigm [67]. SparseBEV reformulates the design by scale-adaptive self-attention and spatio-temporal sampling inspired by AdaMixer[11]. Without explicit dense BEV feature construction, the model excels in detecting foreground objects, with comparable higher translation errors [55]. These sparse architectures cannot be easily applied for full environmental understanding, such as the drivable area and road boundaries [62, 61]. Hence, we focus on dense BEV-based architectures, as they are best suited for the full stack of planning-oriented end-to-end autonomous driving.
2.2 Multi-modal Architectures
Figure 2: Overview of the Height Association Transformer. The radar fusion module exertes a pushing effect into the BEV.
Figure 3: Details of the Radar-weighted Backward Projection. The radar features pull RGB information for refinement.
Initial works on camera-radar fusion have been restricted by the view disparity between the sensor inputs. GrifNet [21], CRAFT [22] and Centerfusion [51] typically follow Region-of-Interest-pooled late-fusion approaches in the perspective or BEV-view. They focus on producing camera-based detections, refined by the offsets to associated radar point detections. Others directly concatenate view-aligned radar feature maps to high-level image features and feed the fused features into a standard 3D detector head [72, 46]. They rely on simple heuristics-based radar point duplication due to the missing height information. We address these shortcomings by designing an end-to-end trainable association to image feature height dimension. X3KD [24] investigates a knowledge distillation framework distilling across modalities from a LiDAR-based teacher, whereas BEVGuide [49] and FUTR3D [5] employ BEV query-based feature sampler with a transformer decoder, but none of them leverages the radar at all for view transformation.
In general, fusing the radar in its natural 3D representation form of the BEV grid has proven to be most effective sofar [78, 12, 23]. RCBEV [85] empirically showed that using specialized or heavy point-cloud-processing feature backbones does not bring performance gains - hence why our state-of-the-artarchitecture leverages efficient pillar-encoding voxelization [25, 29].
The former state-of-the-art in radar-camera fusion, CRN [23] employs the same underlying architecture as BEVDepth [33]. Kim et al. introduce a multi-modal deformable attention mechanism, a BEV-fusion feature operator with a bigger receptive field than simple concatenation [45]. Additionally, they propose a radar-assisted view transformation (RVT) to complement a monocular depth estimation network with a radar-occupancy-based view projection, working in two separate streams. They were the first on nuScenes [2] to show that with an increasing perception range of up to 100m, radar-fusion-based methods outperform LiDAR at far distances [80]. While CRN’s RVT is a step in the right direction, it still relies on two parallel but separate view transformations. We propose a more unified approach to solve this spatial misalignment introduced by monocular depth cues by moving the fusion to a lower level and designing a hybrid fusion.
2.3 Occupancy Prediction
Occupancy Prediction is a new task emerging in the context of 3D perception with the goal of predicting the occupancy and semantic class of every voxel in the surrounding 3D environment [62, 61] It is an important intermediate task in the full stack of planning-oriented end-to-end autonomous driving, positively affecting planning L2 error and collision rate. Semantic occupancy prediction even helps to outperform LiDAR-based planners [15]. Dense architectures (2.1) are well equipped to predict the 3D status of dynamic and static foreground as well as background voxels, as this task is heavily dependent on the quality of depth estimation [35, 53, 81, 48]. Instead of working on the 2D BEV space, the ground plane grid extends to 3D voxels, which are then predicted by 3D convolutional encoders. We show, that a camera-radar fused BEV representation is the best suited for this task, excelling over all previous camera-only methods.
3 HyDRa Architecture
We introduce HyDRa , a novel camera radar fusion architecture aiming to reduce the depth prediction error and leading to state-of-the-art performance in 3D object detection and semantic occupancy prediction. The overall architecture is visualized in Figure 1 and consists of the following key components:
  • Modality-specific feature encoder: Multi-view images are fed into a 2D encoder and converted to high-level feature maps. The radar point cloud is voxelized and encoded to one height level by point pillars [25].
  • Unified Depth Prediction: Our newly proposed Height Association Transformer leverages cross-attention to associate pillar features (missing height, sparse depth) in the respective camera frustrum with image columns (missing depth, dense height). The geometry-aware features are added residually and converted into a dense depth distribution.
  • BEV Fusion: The forward projection module [18] generates the initial BEV representation and is concatenated with the radar pillars channels and fused by a simple Squeeze-and-Exictation layer [14, 37]. Cached history features [55] can be optionally fused.
  • Radar-guided Backward Projection: The backward projection module [34] refines the initial sparse features, guided by the Radar-weighted Depth Consistency calculated from radar BEV features, its implicitly encoded location and radar-aware depth distribution.
  • Downstream Task Heads: The fused BEV representation is encoded by additional residual blocks and fed into the respective task head.
[…]
CenterPoint-P [80] L Pillars - 59.8 49.4 0.320 0.262 0.377 0.334 0.198
CenterPoint-V [80] L Voxel - 65.3 56.9 0.285 0.253 0.323 0.272 0.186
BEVDepth [33] C R50 $256\times 704$ 47.5 35.1 0.639 0.267 0.479 0.428 0.198
FB-BEV [36] C R50 $256\times 704$ 49.8 37.8 0.620 0.273 0.444 0.374 0.200
SOLOFusion [55] C R50 $256\times 704$ 53.4 42.7 0.567 0.274 0.411 0.252 0.188
X3KD [24] C+R R50 $256\times 704$ 53.8 42.3 - - - - -
StreamPETR [67] C R50 $256\times 704$ 54.0 43.2 0.581 0.272 0.413 0.295 0.195
SparseBEV [39] C R50 $256\times 704$ 54.5 43.2 0.606 0.274 0.387 0.251 0.186
CRN [23] C+R R50 $256\times 704$ 56.0 49.0 0.487 0.277 0.542 0.344 0.197
HyDRa C+R R50 $256\times 704$ 58.5 49.4 0.463 0.268 0.478 0.227 0.182
MVFusion [72] C+R R101 $900\times 1600$ 45.5 38.0 0.675 0.258 0.372 0.833 0.196
FUTR3D [5] C+R R101 $900\times 1600$ 50.8 39.9 - - - 0.561 -
BEVDepth [33] C R101 $512\times 1408$ 53.5 41.2 0.565 0.266 0.358 0.331 0.190
BEVGuide [49] C+R R101 $224\times 480$ 53.7 42.1 - - - 0.390 -
SparseBEV [39] C R101 $512\times 1408$ 59.2 50.1 0.562 0.265 0.321 0.243 0.195
StreamPETR [67] C R101 $512\times 1408$ 59.2 50.4 0.569 0.262 0.315 0.257 0.199
CRN [23] C+R R101 $512\times 1408$ 59.2 52.5 0.460 0.273 0.443 0.352 0.180
HyDRa C+R R101 $512\times 1408$ 61.7 53.6 0.416 0.264 0.407 0.231 0.186
Table 1: 3D Object Detection on nuScenes val set. ‘L’, ‘C’, and ‘R’ represent LiDAR, Camera, and Radar, respectively. ( Backbones benefitted from perspective depth pre-training.)
3.1 Vision-Centric Foundation
In our method we propose an extension to existing vision-centric 3D perception systems. Thus we build upon the recent advancements in camera-based high-performance object detection (BEVDet Series [18, 33]) and the simple but effective concept of fusing radar and image features in the BEV space [12]. The setup follows our baseline method CRN [23]. The image encoder, multi-resolution feature pyramid neck [38], the Context-Net, Depth-Net, and BEV-Encoder [13] follow the same architecture choices as previous work [18, 33].
3.2 Height Association Transformer
Our method utilizes the strength of each modality to overcome the main challenge in vision-centric depth prediction. We propose a new transformer-based plug-in module to leverage the complementary radar pillar features (cf. Fig. 3). The primary challenge is associating the radar features $R$ with the respective image features $F$ in the perspective view. For comparison, CRN’s RVT [23] collapses the image features due to missing information in the height dimension and uses the sparse encoded radar-only occupancy in the frustum to splat the features in the BEV space. Instead of using a separate view transformation or the noisy depth map as input [46], we aim to learn which part of the downsampled feature map should benefit from the noisy and ambiguous radar encoding with an extended receptive field along the complete image height.
Let $F\in\mathbb{R}^{B\times N\times H\times W\times C}$ denote 2D image backbone output, while $R\in\mathbb{R}^{B\times N\times 1\times W\times C}$ describes the radar pillar features in the respective camera frustum. $B$ is the batch size, $N$ the number of camera views, $H$ the downsampled feature height dimension, $W$ the downsampled feature width, $D$ the number of depth bins in the camera frustum and $C$ the same-size embedding dimension. For efficiency, we query a single column $w$ of discrete monocular height features $F_{bnw}\in\mathbb{R}^{H\times 1\times C}$ for the $w$-th respective radar feature keys and values $R_{bnw}\in\mathbb{R}^{1\times D\times C}$ on the ground plane. Therefore, we reshape the feature tensors to a new batch size $B^{\prime}=B\times N\times W$. For each ’sequence’ we apply a learnable positional embedding. Next, we encode the height bins with self-attention and decode and fuse the radar features via cross-attention [64]. A peak in activations on the depth bin plane is compared with promising height proposals, encoding the geometric relationship in the attention, inducing sparse but strong metric cues into the dense feature space. Either modality is missing crucial information that the other sensor can provide. To stabilize the learning, we apply a zero-initialized layer norm [1] to the fusion features, and add them residually to the original image features. The resulting feature map $F^{\prime}$ is the input for the depth and context network, producing the final more robust depth distribution.
[…]
PointPillars [25] L Pillars 55.0 40.1 0.392 0.269 0.476 0.270 0.102
CenterPoint [80] L Voxel 67.3 60.3 0.262 0.239 0.361 0.288 0.136
KPConvPillars [63] R Pillars 13.9 4.9 0.823 0.428 0.607 2.081 1.000
CenterFusion [51] C+R DLA34 44.9 32.6 0.631 0.261 0.516 0.614 0.115
MVFusion [72] C+R V2-99 51.7 45.3 0.569 0.269 0.410 0.781 0.203
CRAFT [22] C+R DLA34 52.3 41.1 0.467 0.268 0.456 0.519 0.114
BEVDepth [33] C ConvNeXt-B 60.9 52.0 0.445 0.243 0.352 0.347 0.127
SOLOFusion [55] C ConvNeXt-B 61.9 54.0 0.453 0.257 0.376 0.276 0.148
BEVFormerV2 [75] C InternImage-B 62.0 54.0 0.47.9 0.248 0.320 0.295 0.123
CRN [23] C+R ConvNeXt-B 62.4 57.5 0.416 0.264 0.456 0.365 0.130
StreamPETR [67] C V2-99 63.6 55.0 0.479 0.239 0.317 0.241 0.119
SparseBEV [39] C V2-99 63.6 55.6 0.485 0.244 0.332 0.246 0.117
HyDRa C+R V2-99 64.2 57.4 0.398 0.251 0.423 0.249 0.122
Table 2: 3D Object Detection on the nuScenes [2] test set. By default, V2-99 [54, 26] is pre-trained for depth estimation, whereas ConvNext-B [44] is initialized from ImageNet-22K [6]. The larger ConvNext Base backbone typically leads to better performance [33].
3.3 Radar-Weighted Depth Consistency
In our method we tackle two challenges of spatial misalignments and projection inconsistencies when integrating the depth information from multiple sensors. CRN [23] tries to compensate for spatial misalignments with a global receptive field with their Multi-modal Deformable Cross Attention (MDCA) module, which solely operates within the BEV. Instead, we propose concatenating the initial set of lifted and splatted BEV features with the radar pillar channels and refining them with our radar-weighted back-projection pipeline. Extending the back-projection-as-refinement [36] to the multi-modal domain, we enforce the consistency between the two projection spaces on both sides of the view disparity [85]. We upgrade the full cycle of Depth Consistency three-fold to enhance the projection quality, leveraging the synergy of our HAT module and a lightweight radar guidance network (RGN). With a small $3\times 3$ convolution followed by a sigmoid, the RGN encodes the radar-only BEV features to additional attention weights $r$ (cf. Fig. 3).
The overall concept relies on the classic projection of 3D BEV points $(x,y,z)$ onto 2D image coordinates $(u,v)$ with projection matrix $P$. Every BEV proposal location $Q_{x,y}$ implicitly encodes a depth value $d_{Q}$ for the corresponding camera image projection.
$d\cdot\begin{bmatrix}u&v&1\end{bmatrix}^{T}=P\cdot\begin{bmatrix}x&y&z&1\end{bmatrix}^{T}$ (1)
By weighting the cross-attention [88] with the consistency $w_{C}$ (scalar product) of the predicted depth distribution $d_{P}$ and implicit query depth value $d_{Q}$ (converted to a distribution), we enrich the otherwise sparse, misaligned, and especially unmatched single-modal features.
$w_{c}=d_{P}\cdot d_{Q}$ (2)
Our HAT module tackles the left side of Equation 1 by strengthening the explicit depth (pushing effect [12] of radar). BEV-fusion activates better proposal locations $Q_{x,y}$ for the right side of Equation 1 (pulling effect [12] of radar). The deformable cross-attention brings them closer together and enforces depth consistency on both sides of the equation. We follow the notation of [36] and adapt to the radar-weighted Spatial Cross-Attention (SCA):
$\text{SCA}(Q_{x,y},F)\!=\!\sum\limits_{i=1}^{{N_{\text{c}}}}\sum\limits_{j=1}^{{N_{\text{ref}}}}\mathcal{F}_{d}(Q_{x,y},\mathcal{P}_{i}(x,y,z_{j}),F_{i})\cdot\bm{w_{c}^{ij}}\cdot\bm{r_{x,y}},$
with $\mathcal{F}_{d}$ denoting the deformable attention, dynamically sampling features around the projection point $\mathcal{P}_{i}(x,y,z_{j})$ on the image feature map $F_{i}$, weighted by depth consistency $w_{c}$ and radar attention $r$. Equation 3 highlights where the complementary hybrid fusion approach is beneficial for a more robust depth prediction. Furthermore, we redesigned the ordering of the modules by moving the backprojection after the temporal fusion to capitalize on partially occluded objects indicated not only by radar reflections but also propagated history features.
Methods Input Backbone Image Size mIoU $\uparrow$ Bus Car Ped. Truck MC
MonoScene [3] C R101 $900\times 1600$ 6.1 4.9 9.4 3.0 7.2 4.0
TPVFormer [20] C R101 $900\times 1600$ 27.8 40.8 45.9 18.9 34.2 20.0
BEVFormerOcc [34] C R101 $900\times 1600$ 26.9 40.4 42.4 21.8 30.7 23.9
CTF-Occ [61] C R101 $900\times 1600$ 28.5 38.3 42.2 22.7 31.1 24.5
SparseOcc [40] C R50 $256\times 704$ 30.9 32.9 43.3 23.4 29.3 23.8
BEVStereo-OCC [18] C R50 $256\times 704$ 36.1 42.1 49.6 21.5 37.1 17.4
OctreeOcc [48] C R50 $256\times 704$ 37.4 - - - - -
FlashOcc (M5) [81] C R50 $256\times 704$ 39.0 41.6 50.5 23.8 37.4 23.0
RenderOcc [53] C R50 $256\times 704$ 39.7 - - - - -
FB-OCC [36, 35] C R50 $256\times 704$ 40.7 45.3 50.5 28.2 37.9 29.0
HyDRa C+R R50 $256\times 704$ 44.4 52.3 56.3 35.1 44.1 35.9
Table 3: 3D Occupancy Prediction on Occ3D based on nuScenes val set (CVPR2023 Occupancy Challenge [61] [62]). In addition to the mIoU based on 16 different classes, we also highlight the IoU of the five dynamic classes car, bus, pedestrian (ped.), truck and motorcycle (MC). An overview of all classes (including drivable area and vegetation etc.) can be found in the Appendix.
3.4 Down-stream Tasks
Our final BEV representation can be leveraged for multiple 3D perception tasks, generalizing well for 3D object detection, 3D multi-object tracking (MOT), and 3D semantic occupancy prediction. To decode the features into bounding box characteristics (3D location, 3D size, yaw orientation, and 2D velocity vector), we follow standard practice for dense architectures and use a simple anchor-free center-based head [80]. Similar to [23], we utilize the same box-based tracking-by-detection approach of CenterPoint [80], which relies on a velocity-based greedy distance matching.
Our powerful fused BEV features can be directly converted into a rich semantic occupancy output without the need to forward projecting features into the full 3D voxel cube or a 3D-convolution-based BEV encoder. We use a one-layer 1$\times$1 Channel-To-Height convolution [81] to create the final representation for our occupancy head [35]. We simply enlarge the final BEV features $F\in\mathbb{R}^{B\times X\times Y\times C}$ in the channel dimension $C$ to $C^{\prime}=Z\times C$ and reshape the feature map to the new height dimension $Z$, resulting in $F^{\prime}\in\mathbb{R}^{B\times X\times Y\times Z\times C}$. This new unstacked representation is fed into the low-cost occupancy head [35].
4 Experiments
4.1 Dataset and Metrics
We perform extensive experiments and ablations on HyDRa  using the nuScenes dataset [2] and the Occ3D benchmark [61] and challenge our model with other relevant camera- and radar-based methods on three different tasks.
nuScenes The nuScenes dataset [2] is the main large-scale research benchmark for camera-radar based 3D perception. The dataset contains 1000 diverse and complex scenarios from the Boston and Singapore urban environments, containing left- and right-handed traffic and adverse conditions like night and rain. Annotated at 2Hz, the dataset provides 6 RGB cameras, 5 mmWave radar sensors, and 1 LiDAR sensor. nuScenes provides a rich set of metrics to assess the quality of 3D perception algorithms: mean Average Precision [10], mean Average Translation, Scale, Orientation, Velocity, and Attribute Error. The NDS is a weighted average of all other metrics. The AMOTA score [71] ranks trackers on the benchmark and evaluates the overall 3D multi object tracking performance.
Occ3D The nuScenes Occ3D benchmark [61, 62] evaluates the quality of arbitrarily shaped 3D foreground and background objects. The goal is to predict the complete 3D scene geometry around the ego-car, consisting of a voxelized representation for 18 classes. The benchmark reports the mean Intersection over Union (mIoU) over the number of true positive, false positive, and false negative voxel predictions.
4.2 Implementation Details
Referring to the literature, we train our model in three different scaling settings following the best practices of previous work. We set the default backbones to ResNet50 [13] with an input resolution of 256$\times$704, 118 depth categories, and a pillar and BEV grid size of 0.8m, leading to 128$\times$128 BEV space. We leverage the AdamW [47] optimizer with a batch size of 64. Scaling up the model backbones and image resolutions to ResNet101 [13] (512$\times$1408) and V2-99 [26] (640$\times$1600), the BEV space is doubled to a standard of 256$\times$256. For comparison to the latest camera-based methods on the test set, we opt for the smaller and more efficient V2-99 backbone.
We build upon the code base of the BEVDet Series [18] and in particular employ the efficient implementation of BEVPoolV2 [17] as a forward projection view transformer. We train for 20 epochs following the memory-efficient sequential sampling and settings of [55] without CBGS [87]. This reduces the step-to-step sampling diversity but speeds up the training time significantly. Regarding the backward projection module, we follow the configuration of FB-BEV [36]. To trade off the number of parameters, we half the backbone output and depth net middle channels. Our HAT module is implemented by one layer of vanilla self- and cross-attention working on 16-times downsampled height queries. This leads to sequences of 16, 32, and 40 height bins for the three model sizes. For temporal fusion, we follow [36, 39, 67] and use the last eight history frames. Data augmentations are the same as BEVDet [18] and CRN [23], but no test-time augmentations are used. For a fair evaluation, each model is trained only on the respective task and leverages the same task head and supervision signals as the baseline methods.
[…]
CenterPoint [80] L Voxel 63.8 0.555 18612 22928 760
UVTR [30] C V2-99 51.9 1.125 14994 39209 2204
ByteTrackV2 [83] C V2-99 56.4 1.005 18939 33531 704
StreamPETR [76] C ConvNeXt-B 56.6 0.975 21268 31484 784
CRN [23] C+R ConvNeXt-B 56.9 0.809 16822 41093 946
HyDRa C+R V2-99 58.4 0.950 13996 32950 791
Table 4: 3D Object Tracking on nuScenes test set.
4.3 Main Results
Figure 4: Qualitative comparison of the semantic occupancy prediction in a challenging night scenario. The top row shows the front-view input cameras. We compare FB-OCC [35] with our proposed HyDRa . While the baseline struggles to distinguish different objects at distance, HyDRa showcases spatial consistency and robustness of the detected cars (orange), truck (red) and pedestrian (blue).
3D Object Detection. In Table 1 we report the results of our HyDRa  model on the nuScenes validation set. With a ResNet50 backbone, we achieve a new state-of-the-art of 58.5 NDS, surpassing the previous best score of 56.0 NDS by CRN [23] by 2.5 NDS. Scaling the model and image resolution to ResNet101 and 512$\times$1408, we maintain the large margin and increase the NDS to 61.7, outperforming all other camera- and radar-based methods on the val split. We want to highlight the impressive improvement in the mean Average Translation Error (mATE) of 0.416, which can be attributed to our model’s robust depth estimation concept. Our proposed HyDRa  surpasses CRN with the more powerful ConvNext-Base[33] backbone by 1.4 NDS due to strengthened translation and velocity estimation, leading to a new state-of-the-art of 64.2 NDS on the nuScenes test set (see Table 2). Utilizing a bigger backbone, e.g., Vit-L [9, 67], our method can surpass voxel-based CenterPoint [80]. Qualitative results are published in the Appendix.
Additionally, we show in Table 5 that HyDRa generalizes on the recent View-of-Delft validation set [52], utilizing a more advanced 4D millimeter-wave radar. HyDRa outperforms the best method by an increase of +4.57 mAP in the entire annotation area and +4.06 mAP in the region of interest (compared to the CRN-inspired LXL model [73]).
3D Multi Object Tracking. We summarize the tracking results in Table 4. Benefitting from accurate object localization and high-quality velocity estimation, the center-based tracking performance of HyDRa  is competitive with the state-of-the-art methods, leading to the best trade-off between false positives (FP), false negatives (FN), and identity switches (IDS) on the nuScenes test set, once more narrowing the gap to LiDAR-based CenterPoint.
Methods Input AP in the Entire Annotated Area (%) AP in the Region of Interest (%)
Car Pedestrian Cyclist mAP Car Pedestrian Cyclist mAP
PointPillar [25] R 37.06 35.04 63.44 45.18 70.15 47.22 85.07 67.48
RadarPillarNet [84] R 39.30 35.10 63.63 46.01 71.65 42.80 83.14 65.86
RCFusion [84] C+R 41.70 38.95 68.31 49.65 71.87 47.50 88.33 69.23
SMURF [79] R 42.31 39.09 71.50 50.97 71.74 50.54 86.87 69.72
LXL [73] C+R 42.33 49.48 77.12 56.31 72.18 58.30 88.31 72.93
HyDRa C+R 52.83 56.57 73.25 60.88 80.65 62.90 87.43 76.99
Table 5: 3D object detection results on View-of-Delft  [52]. The RoI ($(x,y,z)|-4m<x<4m,z<25m$) is the driving corridor close to the ego-vehicle. The IoU thresholds are set to 0.5 for cars and 0.25 for pedestrians and cyclists (official evaluation settings).
3D Occupancy Prediction.
Dense architectures with strong depth prediction excel in occupancy prediction [61, 62]. Our BEV-based architecture showcases superior performance on the Occ3D benchmark, outperforming all other camera-based methods by a large margin of 3.7 mIoU (see Table 3). It works remarkably well on challenging dynamic objects as visualized in Figure Fig. 4. Complementary camera-radar fusion is the key to unlocking the potential of vision-centric 3D sensing. While sparse representations perform well in detection (cf. SparseBEV [39] in Table 2), they underperform in occupancy prediction (cf. Sparseocc [40] in Table 3). This emphasizes the importance of a dense representation for full 3D scene understanding.
4.4 Ablation Studies
[…]
BEVDepth [33] C 44.4 32.1 0.737 0.344
CRN RVT [23] C+R 45.6 38.3 0.619 0.510
BEVDepth + RF C+R 48.1 37.2 0.614 0.284
BEVDepth + HAT C+R 50.3 39.8 0.586 0.287
CRN [23] C+R 48.9 39.7 0.616 0.361
RDC C+R 52.7 40.5 0.534 0.261
HyDRa C+R 53.6 41.2 0.493 0.257
Table 6: Ablation of fusion strategies: RF - Baseline Range-Fusion of projected radar depth maps; CRN with occupancy-based Radar View Transformation (RVT); HyDRa  with Radar-weighted Depth Consistency (RDC) and Height Association Transformer (HAT) is our full model. We train and evaluate in the front-view setting.
To validate the individual contribution of each of our proposed designs and components, we conducted detailed ablation studies. The ablation studies and baseline training of Tab. 6, Tab. 7, and Tab. 8 are performed in the front-view setting. Input for the model are the three forward-facing cameras and the corresponding radar sensors, which reduces the data and computation size by a factor of two. The ablations are conducted on the nuScenes validation set.
Effect of the Range View Fusion. We develop a simple baseline similar to Radiant [46] to show that using radar in a unified depth net brings more performance than separate branches. We extend BEVDepth [33] with an additional ResNet18 [13] encoding the radar in the same image view. By projecting radar points into the image plane, we create a range image of the radar point cloud, converting 3D points into a pixel location. Instead of RGB channels, radar channels are used, with every non-radar pixel set to zero. This pseudo-image is input to a smaller ResNet18 encoder, and each output stage is concatenated to the respective RGB encoder output.
We show in Table 6 that this unified approach already outperforms the RVT of CRN. We indicate, that the velocity estimation is also benefitting from a perspective fusion, compared to pure radar-occupancy-guided transformation of the semantic features. Processing the mostly sparse stripe of radar points with 2D convolutions is neither effective nor efficient. To overcome the limitations of the naive baseline, we propose a more sophisticated approach, the HAT module, which is a novel way of fusing radar and camera features in the perspective view and is capable of significantly improving the depth estimation of camera-based models. Compared to related work, we do not have to rely on heuristic height extensions [46, 51] or neglect it at all [12] by collapsing image features [23].
Effect of Radar Weighted Depth Consistency. Radar-informed depth consistency is a crucial concept of our architecture, showing significant and consistent improvements over the baseline architecture of CRN. Especially adding the HAT module into this paradigm leverages synergistic effects, leading to strong metric depth understanding (cf. Tab. 6). Without temporal or depth supervision, as ablated in Table 7, HyDRa  is the superior concept for increased depth and velocity sensing, showing a decent improvement over the camera-only baseline. This could be important for time-critical situations e.g. collision avoidance [78].
Order Matters. In Table 8, we show that the order of the fusion matters. The radar BEV features should be fused before any 4D aggregation and backward projection. Enriching the queries with radar and temporal information, poses more meaningful cues for refinement, without additional computational cost as the modules are working on the same BEV channel dimensions.
[…]
BEVDepth [33] C 33.3 26.7 0.803 0.930
CRN [23] C+R 42.5 35.0 0.607 0.584
HyDRa C+R 45.8 34.7 0.569 0.482
Table 7: Comparison of BEVDepth, CRN and HyDRa without temporal fusion or depth supervision. We train and evaluate in the front-view setting.
[…]
A BP $\rightarrow$ 4D $\rightarrow$ RF C+R 52.0 41.1 0.551 0.261
B RF $\rightarrow$ BP $\rightarrow$ 4D C+R 52.7 41.2 0.527 0.281
C RF $\rightarrow$ 4D $\rightarrow$ BP C+R 53.6 41.2 0.493 0.257
Table 8: Ablation of component order (BP: Backprojection, 4D: Temporal Fusion, RF: Radar BEV-Fusion). Radar Attention for Backprojection is not available in setting A. This reports our full model, including HAT and RDC. We train and evaluate in the front-view setting.
4.5 Limitations
In future endeavors, we want to explore the potential of object-level motion modeling [67] within our dense paradigm, as the ego-motion compensation [16] does not account for highly dynamic objects. Moreover, the orientation poses a bottleneck as sparse radar points only give a rough estimate of the object’s location. Therefore, 4D imaging radars pose an interesting research opportunity [52, 7].
5 Conclusion
We introduce HyDRa, the new state-of-the-art hybrid fusion paradigm, excelling in various 3D perception tasks and showing a promising path for future research in the field of radar-based detection. HyDRa  beats the former state-of-the-art in camera-radar fusion by a clear margin of 1.8 NDS and naturally generalizes to the first radar-enhanced occupancy prediction model, improving upon the best camera baseline by 3.7 mIoU. With this line of work, we want to contribute to safer autonomous driving due to better handling of low-visible objects, more robust depth estimation, and accurate velocity estimation.